A Bibliography of Publications in American Mathematical Monthly: 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 January 2012 Version 1.09

Title word cross-reference

#12355 [138]. #22428 [28].

(0, 1) [112]. (0, ∞) [112]. (_{k}^{m}) [59]. a^{r+s} ≡ a^{r} (mod m) [138, 48]. k [1]. k + 1 [1]. m(m − 1) ⋯ (m − h + 1) [59]. n [78]. π [121]. R^{n} [110]. V [92]. y′ = sin(xy) [123].

-blocks [1]. -flexing [92].

01019a [139].

57 [138]. 58 [28].

80e [140, 141]. 80f [139]. 85 [138, 67, 140, 28, 141, 139]. abstract [26]. according [75]. addendum [139]. again [32]. Airy [76]. algebra [93, 130]. algebraic [17]. algebras [29]. algorithm [121, 133]. alternating [113]. always [129]. Amer [138, 67, 140, 28, 141, 139]. analysis [22, 64]. Analytic [21]. analyticity [16, 100]. antagonism [74]. anti [58]. anti-plane [58]. application [81]. applications [73, 114]. applied [74]. Apportionment [44]. approximate [135]. Approximately [29]. Approximation [26]. arbitrarily [24]. arithmetic [109]. asymptotic [39]. automorphisms [118]. Award [54, 55]. axiomatic [41].

ball [106, 21]. Banach [63]. Banach-Tarski [63]. bases [15]. basis [71]. Begle [25]. behavior [39]. billiards [69]. binomial [45, 5, 6]. blocks [1]. Bonnesen [50]. Bonnesen-style [50]. Borsuk [127]. Borsuk-Ulam [127]. boundaries [26]. boundary [64]. bounded [105]. bounds [72]. Brouwer [131, 21].

calculation [121]. came [66]. cannot [43]. Cantor [104]. Cantor-Lebesgue [104]. cardinality [137]. case [123, 40]. Cauchy [32, 90]. changing [133]. characterization [38, 79, 17]. characterizations [16, 100]. Chauvenet [55]. circle [47]. circles [79]. circular [118]. classical [81]. coefficients [96, 5, 6]. Comment [138]. Comments [134, 97]. commutators [129]. compact [9, 12]. complements [134]. Complex [53, 35]. Computer [4, 74, 123]. concept [34, 139]. conditioning [135]. Conformal [118]. congruence [138, 48]. conics [91]. connected [30, 34]. connected’ [139]. consistency [105]. contain [79]. continuity [125, 132]. continuous [135, 13]. continuously [60]. convergence [116]. convergent [31]. Convex [43, 61]. Correction [139]. Corrections [67, 140, 28]. Coupling [46]. creation [130]. curve [91]. cyclic [7, 11]. Cyclotomic [36, 140, 141].

days [85]. decomposition [96, 51]. density [83]. Dependent [80]. derivatives [39]. difference [117]. differentiable [110, 90]. differential [120]. differentiation [62]. diffusion [64]. digit [87]. Digital [111]. dilemma [56]. dimensional [83]. diphages [10, 19]. Discovering [123]. Dissections [30]. distinguished [54]. distributed [47]. Divisibility [59]. division [27]. domain [30]. doubly [61]. Dr [55]. drafts [56].

E2307 [1]. E2315 [2]. E2328 [3]. E2492 [5, 6]. E2613 [9, 12]. E2636 [10, 19]. Edward [25]. eigenvalues [35].

Elementary [5, 6, 9, 10, 12, 19, 65, 82, 83]. equation [32]. equations [120]. Ericksen [58]. error [72]. Euclid [75]. Euclidean [125]. Euler [86, 93]. Evolution [34, 139, 22]. example [129]. existence [52]. exponential [122]. Extended [14, 28]. extensions [17]. Extrema [128]. Extremal [20].

factorization [36, 140, 141]. Fermat [40]. field [38]. Finding [112]. finite [38, 58, 29]. first [66]. fixed [131, 21]. fixed-point [131, 21]. flexing [92]. flow [64]. forms [135]. Fourier [66]. fourteenth [18]. fraction [96]. function [97, 13, 78, 61]. functional [32]. Functions [24, 68, 51, 128, 135]. Further [97].

game [77, 131]. generalization [68, 88]. generalized [99]. generators [7, 11]. Geometrical [91]. geometries [125]. Geometry [75, 41]. given [87]. Gödel [124]. Golomb’s [141]. Gradient [16, 100]. graphs [89]. Grassmann [130]. Griffith [25]. Groups [3, 53].

hairy [106, 21]. harmonic [22, 31]. heat [64]. Hermann [130]. Hex [131]. hexahexaflexagon [92]. Hilbert [18]. homotopy [90]. House [44].

identities [23]. identity [82].

improvement [45]. incompleteness [124]. increasing [13]. independent [80]. inequalities [70, 94, 50]. inequality [81, 78]. infinite [107]. inseparable [17]. integer [95, 119]. integers [87, 67]. intervals [77]. introduction [57]. Irregular [67]. isoperimetric [94, 50]. iterate [78].

J [55]. Jordan [41].

Kantorovich [72]. Kronecker [42].

L [138]. Laguerre [114]. last [40]. lattice [126]. laws [49]. Lebesgue [104]. lemma [120]. length [124]. length-of-proof [124]. linear [127, 130, 120]. lines [27]. Livingston [138, 138]. Local [135]. lost [57].

M [138, 67]. manifolds [84]. manipulative [65]. many [112]. Math [138, 67, 140, 28, 141, 139]. Mathematics [102, 4, 81, 73, 103]. Matrices [35, 61]. maximization [8, 33]. mean [14, 28]. measurable [52]. Measure [137]. metacyclic [53]. method [65]. methods [44]. metric [110]. miniature [58]. Minimal [107, 133]. minimal-path [133]. missing [87]. mnemonic [65]. mods [5, 6]. money [133]. Monthly [138, 67, 140, 28, 141, 139]. Morley [37]. Moser [94]. moving [64]. MR [138, 140, 28, 141, 139].

Neil [55]. Neugebauer [54]. Neumann [29]. no [138, 67, 140, 28, 141, 110, 139]. non [52, 125]. non-Euclidean [125]. non-measurable [52]. noncommutative [22]. Nonstandard [115]. normal [71]. normed [127]. note [95, 138, 141]. notebook [57].

objective [102]. obstreperous [86]. officer [93]. old [114]. one [83]. one-dimensional [83]. optimal [72]. optimization [101]. order [120]. orthogonality [127]. Otto [54].

paradox [63]. parametric [135]. partial [96]. partitions [95]. path [133]. penosculating [91]. perception [51]. Perfect [136]. periodic [51]. periods [24]. permanent [61]. permutation [38, 7, 11]. perspective [124]. Picturesque [122]. Pixley [60]. Pixley-Roy [60]. plane [132, 58, 43, 91, 27, 30]. point [131, 21]. points [79, 126]. polygons [43]. polynomial [112]. Polynomials [47, 117, 20, 38, 36, 140, 108, 141]. population [81]. precisely [1]. Prismatoid [99]. prismoid [99]. Prisoner [56]. Prize [55]. Problem [3, 1, 5, 7, 8, 6, 9, 10, 12, 11, 19, 33, 93, 18, 58, 2, 94, 133, 64, 126]. problems [20]. product [23, 70]. product-sum [23]. Products [129, 89]. professional [56]. progressions [109]. projective [35, 132]. Proof [40, 98, 124, 106, 42, 90, 82, 83]. proofs [21]. properties [91]. property [108]. prototype [64]. pure [73].

quadratic [108]. quadrilaterals [136]. questions [109]. quotients [117].

Ramanujan [57]. random [8, 33, 80]. Rational [49, 79]. real [132]. reciprocals [87]. reciprocity [49]. regions [118]. regulated [68]. related [94]. Relation [4]. Remarks [69]. renewal [46]. representations [53]. Representatives [44]. role [73]. Rolle [98]. roots [112]. Roy [60].

S. [141]. Schauder [15]. Schoenflies [41]. Schreiber [67]. Schröder [2]. Science [4, 102, 74]. Search [101]. second [120]. second-order [120]. semimetrizable [60]. sequence [116]. Sequences [1]. series [86, 66, 45, 113, 31]. service [54]. set [115, 61]. sets [9, 12, 52]. several [128]. shear [58]. sides [95, 119]. Silverman [77]. simple [45, 42]. simply [30]. simply-connected [30]. singular [13]. Sloane [55]. small [24]. Solution [1, 6, 12, 11, 33, 3, 2]. solutions [78]. Some [23, 31, 109, 114]. space [35, 60]. spaces [127, 107]. special [40]. splitting [17]. sports [56]. stability [64]. states [58]. Steinhaus [126]. stochastic [61]. strong [62]. Student [85]. style [50]. subdues [86]. subseries [31]. subsets [80]. sum [23].

Summing [113]. Sums [87, 122]. survey [65]. symbiosis [74]. Symmetric [62].

Tarski [63]. Taylor [117]. tests [116]. th [78]. their [89]. theorem [98, 46, 127, 104, 124, 106, 131, 42, 72, 21, 88, 37, 40, 105, 90, 71, 83]. theorems [36, 140, 41, 141, 123]. Theoretical [74]. theory [115, 120]. There [110]. tile [43]. Tone [51]. topological [107, 34, 139]. topology [111]. transform [76]. transpositions [89]. treatment [70]. Triangles [119, 95, 84]. trigonometric [82]. triphages [10, 19]. trisector [37]. type [70].

Ulam [127]. ultimate [98]. umbral [65]. Unified [70]. uniformly [47]. unit [47]. Universal [93, 58]. use [90]. used [135]. useful [120]. uses [65].

values [14, 28]. variables [128, 80]. variation [97]. very [86].

W [141]. Weierstrass [70]. which [79]. without [125, 132]. work [114].

zeros [47]. Zolotarev [88].

References

Knuth:1972:SPK

[1] Donald E. Knuth. Sequences with precisely k + 1 k-blocks: Solution to problem E2307. American Mathematical Monthly, 79(7):773–774, September 1972. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1972:SPS

[2] Donald E. Knuth. Schröder’s problem: Solution to problem E2315. American Mathematical Monthly, 79(8):910, October 1972. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1972:GSP

[3] Donald E. Knuth. Groups: Solution to problem E2328. American Mathematical Monthly, 79(10):1138–1139, December 1972. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1974:CSR

[4] Donald E. Knuth. Computer science and its relation to mathematics. American Mathematical Monthly, 81(??):323– 343, April 1974. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). A shorter form of this article appeared in American Scientist 61 (1973), 707–713; reprinted in Computers and People 23,9 (September 1974), 8– 11; and in Mathematics: People, Problems, Results, ed. by Douglas M. Campbell and John C. Higgins, vol. 3 (Belmont, Calif.: Wadsworth, 1984), 37– 47. Hungarian translation in Matematikai Lapok 24 (1973, published 1975), 345–363. Slovenian translation in Obzornik za Matematiko in Fiziko 22 (1975), 129–138, 161–167. Slovak translation (abridged) in Pokroky Matematiky, Fiziky a Astronomie 21 (1976), 88–96. Russian translation by Natal’�ıa G. Gurevich in Sovremennye Problemy Matematiki 11,12 (Moscow: Znanie, 1977), 4–32.

Knuth:1974:EPE

[5] Donald E. Knuth. Elementary problem E2492 (binomial coefficients and mods). American Mathematical Monthly, 81: 902, 1974. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [6].

Knuth:1975:SEP

[6] Donald E. Knuth. Solution to elementary problem E2492 (binomial coefficients and mods). American Mathematical Monthly, 82:855, 1975. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [5].

Knuth:1975:PCP

[7] Donald E. Knuth. Problem 6049 (cyclic permutation generators). American Mathematical Monthly, 82:856, 1975. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [11].

Knuth:1975:PRM

[8] Donald E. Knuth. Problem 6050 (random maximization). American Mathematical Monthly, 82:856, 1975. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [33].

Knuth:1976:EPE

[9] Donald E. Knuth. Elementary problem E2613 (compact sets). American Mathematical Monthly, 83:656, 1976. ISSN 0002-9890 (print), 1930-0972 (electronic). Solution in [12].

Knuth:1977:EPE

[10] Donald E. Knuth. Elementary problem E2636 (diphages and triphages). American Mathematical Monthly, 84: 134, 1977. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1977:SPCa

[11] Donald E. Knuth. Solution to problem 6049 (cyclic permutation generators). American Mathematical Monthly, 84:397, 1977. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [7].

Knuth:1977:SEP

[12] Donald E. Knuth. Solution to elementary problem E2613 (compact sets). American Mathematical Monthly, 84: 827–828, 1977. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [9].

Takacs:1978:ICS

[13] Lajos Takács. An increasing continuous singular function. American Mathematical Monthly, 85(1):35–37, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:1978:EMV

[14] Anonymous. Extended mean values. American Mathematical Monthly, 85(2): 84–90, ???? 1978. See corrections [28].

Halmos:1978:SB

[15] P. R. Halmos. Schauder bases. American Mathematical Monthly, 85(4):256– 257, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bear:1978:GCA

[16] H. S. Bear and G. N. Hile. Gradient characterizations of analyticity. American Mathematical Monthly, 85(5):333– 337, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Norris:1978:CSI

[17] Michael J. Norris and William Yslas Vélez. A characterization of the splitting of inseparable algebraic extensions.

American Mathematical Monthly, 85(5): 338–341, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Humphreys:1978:HFP

[18] J. E. Humphreys. Hilbert’s fourteenth problem. American Mathematical Monthly, 85(5):341–353, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1978:EPE

[19] Donald E. Knuth. Elementary problem E2636 (diphages and triphages). American Mathematical Monthly, 85:385–386, 1978. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boas:1978:EPP

[20] R. P. Boas. Extremal problems for polynomials. American Mathematical Monthly, 85(6):473–475, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Milnor:1978:APH

[21] John Milnor. Analytic proofs of the “hairy ball theorem” and the Brouwer fixed-point theorem. American Mathematical Monthly, 85(7):521–524, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gross:1978:ENH

[22] Kenneth I. Gross. On the evolution of noncommutative harmonic analysis. American Mathematical Monthly, 85(7): 525–548, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Carlitz:1978:SPS

[23] L. Carlitz. Some product-sum identities. American Mathematical Monthly, 85(7):570–572, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cignoli:1978:FAS

[24] R. Cignoli and J. Hounie. Functions with arbitrarily small periods. American Mathematical Monthly, 85(7):582– 584, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zelinka:1978:EGB

[25] Martha Zelinka. Edward Griffith Begle. American Mathematical Monthly, 85 (8):629–631, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bauer:1978:AAB

[26] Heinz Bauer. Approximation and abstract boundaries. American Mathematical Monthly, 85(8):632–647, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wetzel:1978:DPL

[27] John E. Wetzel. On the division of the plane by lines. American Mathematical Monthly, 85(8):647–656, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leach:1978:CEM

[28] E. B. Leach and M. C. Sholander. Corrections to: “Extended mean values” [Amer. Math. Monthly 85 (1978), no. 2, 84–90, MR 58 #22428]. American Mathematical Monthly, 85(8):656, ????

1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [14].

Moore:1978:AFN

[29] Calvin C. Moore. Approximately finite von Neumann algebras. American Mathematical Monthly, 85(8):657–659, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wetzel:1978:DSC

[30] John E. Wetzel. Dissections of a simply-connected plane domain. American Mathematical Monthly, 85(8):660– 661, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wadhwa:1978:SCS

[31] A. D. Wadhwa. Some convergent subseries of the harmonic series. American Mathematical Monthly, 85(8):661– 663, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Letac:1978:CFE

[32] Gérard Letac. Cauchy functional equation again. American Mathematical Monthly, 85(8):663–664, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1978:SPR

[33] Donald E. Knuth. Solution to problem 6050 (random maximization). American Mathematical Monthly, 85:686–688, 1978. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [8].

Wilder:1978:ETC

[34] R. L. Wilder. Evolution of the topological concept of “connected”. American Mathematical Monthly, 85(9):720– 726, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [139].

Alexander:1978:MEC

[35] J. C. Alexander. Matrices, eigenvalues, and complex projective space. American Mathematical Monthly, 85(9):727– 733, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Golomb:1978:CPF

[36] Solomon W. Golomb. Cyclotomic polynomials and factorization theorems. American Mathematical Monthly, 85(9): 734–737, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See corrections [140] and note [141].

Oakley:1978:MTT

[37] Cletus O. Oakley and Justine C. Baker. The Morley trisector theorem. American Mathematical Monthly, 85(9):737– 745, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Carlitz:1978:CPP

[38] L. Carlitz and Jo Ann Lutz. A characterization of permutation polynomials over a finite field. American Mathematical Monthly, 85(9):746–748, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boas:1978:ABD

[39] R. P. Boas, H. Pollard, and D. V. Widder. The asymptotic behavior of derivatives. American Mathematical Monthly, 85(9):749–750, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Powell:1978:PSC

[40] Barry Powell. Proof of a special case of Fermat’s last theorem. American Mathematical Monthly, 85(9):750–751, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guggenheimer:1978:JST

[41] H. Guggenheimer. The Jordan and Schoenflies theorems in axiomatic geometry. American Mathematical Monthly, 85(9):753–756, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Greiter:1978:SPT

[42] Gebhard Greiter. A simple proof for a theorem of Kronecker. American Mathematical Monthly, 85(9):756–757, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Niven:1978:CPC

[43] Ivan Niven. Convex polygons that cannot tile the plane. American Mathematical Monthly, 85(10):785–792, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Saari:1978:AMH

[44] Donald G. Saari. Apportionment methods and the House of Representatives. American Mathematical Monthly, 85 (10):792–802, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gerber:1978:SIB

[45] Leon Gerber. A simple improvement on the binomial series. American Mathematical Monthly, 85(10):808–809, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Athreya:1978:CRT

[46] K. Athreya, D. McDonald, and P. Ney. Coupling and the renewal theorem. American Mathematical Monthly, 85 (10):809–814, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

OHara:1978:PZU

[47] Patrick J. O’Hara and Rene S. Rodriguez. Polynomials with zeros uniformly distributed on the unit circle. American Mathematical Monthly, 85 (10):814–817, ???? 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Livingston:1978:C

[48] A. E. Livingston and M. L. Livingston. The congruence a^{r+s} ≡ a^{r} (mod m). American Mathematical Monthly, 85(2): 97–100, ???? 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See comment [138].

Lehmer:1978:RRL

[49] Emma Lehmer. Rational reciprocity laws. American Mathematical Monthly, 85(??):467–472, ???? 1978.

Osserman:1979:BSI

[50] Robert Osserman. Bonnesen-style isoperimetric inequalities. American Mathematical Monthly, 86(1):1–29, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gale:1979:TPD

[51] David Gale. Tone perception and decomposition of periodic functions. American Mathematical Monthly, 86(1): 36–42, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mauldin:1979:ENM

[52] R. Daniel Mauldin. The existence of non-measurable sets. American Mathematical Monthly, 86(1):45–46, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Basmaji:1979:CRM

[53] B. G. Basmaji. Complex representations of metacyclic groups. American Mathematical Monthly, 86(1):47– 48, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boas:1979:ADS

[54] R. P. Boas. Award for distinguished service to Otto Neugebauer. American Mathematical Monthly, 86(2):77– 78, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Roselle:1979:ACP

[55] David P. Roselle. Award of the Chauvenet Prize to Dr. Neil J. A. Sloane. American Mathematical Monthly, 86(2):

79, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brams:1979:PDP

[56] Steven J. Brams and Philip D. Straffin, Jr. Prisoner’s dilemma and professional sports drafts. American Mathematical Monthly, 86(2):80–88, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andrews:1979:IRL

[57] George E. Andrews. An introduction to Ramanujan’s “lost” notebook. American Mathematical Monthly, 86(2):89– 108, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knowles:1979:USF

[58] James K. Knowles. Universal states of finite anti-plane shear: Ericksen’s problem in miniature. American Mathematical Monthly, 86(2):109–113, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Harborth:1979:D

[59] Heiko Harborth. Divisibility of (_{k}^{m}) by m(m−1) ⋯ (m−h+1). American Mathematical Monthly, 86(2):115–117, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kenton:1979:PRS

[60] S. A. Kenton and S. D. Shore. The Pixley-Roy space is continuously semimetrizable. American Mathematical Monthly, 86(2):117–119, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wang:1979:WPF

[61] Edward T. H. Wang. When is the permanent function convex on the set of doubly stochastic matrices. American Mathematical Monthly, 86(2):119– 121, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Belna:1979:SSD

[62] C. L. Belna, M. J. Evans, and P. D. Humke. Symmetric and strong differentiation. American Mathematical Monthly, 86(2):121–123, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stromberg:1979:BTP

[63] Karl Stromberg. The Banach-Tarski paradox. American Mathematical Monthly, 86(3):151–161, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Oulton:1979:SAP

[64] D. B. Oulton, D. J. Wollkind, and R. N. Maurer. A stability analysis of a prototype moving boundary problem in heat flow and diffusion. American Mathematical Monthly, 86(3):175–186, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guinand:1979:UMS

[65] Andrew P. Guinand. The umbral method: a survey of elementary mnemonic and manipulative uses. American Mathematical Monthly, 86(3): 187–195, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bochner:1979:FSC

[66] Salomon Bochner. Fourier series came first. American Mathematical Monthly, 86(3):197–199, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Garten:1979:CII

[67] Scott Garten. Corrections to: “Irregular integers” (Amer. Math. Monthly 85 (1978), no. 3, 165–172) by M. Schreiber. American Mathematical Monthly, 86(3): 199, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Davison:1979:GRF

[68] T. M. K. Davison. A generalization of regulated functions. American Mathematical Monthly, 86(3):202–204, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sine:1979:RB

[69] Robert Sine and Vladislav Kreĭnovič. Remarks on billiards. American Mathematical Monthly, 86(3):204–206, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

El-Neweihi:1979:UTI

[70] Emad El-Neweihi and Frank Proschan. Unified treatment of inequalities of the Weierstrass product type. American Mathematical Monthly, 86(3):206–208, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Waterhouse:1979:NBT

[71] William C. Waterhouse. The normal basis theorem. American Mathematical Monthly, 86(3):212, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Miel:1979:KTO

[72] G. J. Miel. The Kantorovich theorem with optimal error bounds. American Mathematical Monthly, 86(3):212– 215, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bernstein:1979:RAP

[73] Dorothy L. Bernstein. The role of applications in pure mathematics. American Mathematical Monthly, 86(4):245– 253, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Harrow:1979:TAC

[74] Keith Harrow. Theoretical and applied computer science: antagonism or symbiosis. American Mathematical Monthly, 86(4):253–260, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fisher:1979:GAE

[75] J. C. Fisher. Geometry according to Euclid. American Mathematical Monthly, 86(4):260–270, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Widder:1979:AT

[76] D. V. Widder. The Airy transform. American Mathematical Monthly, 86(4): 271–277, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Evans:1979:SGI

[77] Ronald J. Evans. Silverman’s game on intervals. American Mathematical Monthly, 86(4):277–281, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Turdza:1979:SII

[78] Erwin Turdza. The solutions of an inequality for the nth iterate of a function. American Mathematical Monthly, 86(4):281–283, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Humke:1979:CCW

[79] Paul D. Humke and Lawrence L. Krajewski. A characterization of circles which contain rational points. American Mathematical Monthly, 86(4):287– 290, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wang:1979:DRV

[80] Y. H. Wang. Dependent random variables with independent subsets. American Mathematical Monthly, 86(4):290– 292, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anderson:1979:ACI

[81] David H. Anderson. The application of a classical inequality to population mathematics. American Mathematical Monthly, 86(4):294–296, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Young:1979:EPT

[82] Robert M. Young. An elementary proof of a trigonometric identity. American Mathematical Monthly, 86(4):296– 297, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zajicek:1979:EPO

[83] Luděk Zajíček. An elementary proof of the one-dimensional density theorem. American Mathematical Monthly, 86(4): 297–298, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chern:1979:TM

[84] Shiing Shen Chern. From triangles to manifolds. American Mathematical Monthly, 86(5):339–349, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brenner:1979:SD

[85] J. L. Brenner. Student days—1930. American Mathematical Monthly, 86(5): 350–356, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Barbeau:1979:ESV

[86] E. J. Barbeau. Euler subdues a very obstreperous series. American Mathematical Monthly, 86(5):356–372, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Baillie:1979:SRI

[87] Robert Baillie. Sums of reciprocals of integers missing a given digit. American Mathematical Monthly, 86(5):372– 374, ???? 1979. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic).

Morton:1979:GZT

[88] Patrick Morton. A generalization of Zolotarev’s theorem. American Mathematical Monthly, 86(5):374–375, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Higgs:1979:PTT

[89] Denis Higgs and Paul de Witte. On products of transpositions and their graphs. American Mathematical Monthly, 86(5):376–380, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Vyborny:1979:UDH

[90] R. Výborný. On the use of a differentiable homotopy in the proof of the Cauchy theorem. American Mathematical Monthly, 86(5):380–382, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schot:1979:GPP

[91] Steven H. Schot. Geometrical properties of the penosculating conics of a plane curve. American Mathematical Monthly, 86(6):449–457, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McLean:1979:FH

[92] T. Bruce McLean. V -flexing the hexahexaflexagon. American Mathematical Monthly, 86(6):457–466, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Evans:1979:UAE

[93] Trevor Evans. Universal algebra and Euler’s officer problem. American Mathematical Monthly, 86(6):466–473, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lutwak:1979:IIR

[94] E. Lutwak. On isoperimetric inequalities related to a problem of Moser. American Mathematical Monthly, 86(6):476– 477, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andrews:1979:NPT

[95] George E. Andrews. A note on partitions and triangles with integer sides. American Mathematical Monthly, 86(6):477– 478, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eustice:1979:CPF

[96] Dan Eustice and M. S. Klamkin. On the coefficients of a partial fraction decomposition. American Mathematical Monthly, 86(6):478–480, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Russell:1979:FCV

[97] A. M. Russell. Further comments on the variation function. American Mathematical Monthly, 86(6):480–482, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Abian:1979:UPR

[98] Alexander Abian. An ultimate proof of Rolle’s theorem. American Mathematical Monthly, 86(6):484–485, ???? 1979.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bradley:1979:PPG

[99] A. Day Bradley. Prismatoid, prismoid, generalized prismoid. American Mathematical Monthly, 86(6):486–490, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Walsh:1979:MGC

[100] Bertram Walsh. More on gradient characterizations of analyticity. American Mathematical Monthly, 86(6):490– 491, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Koopman:1979:SO

[101] Bernard O. Koopman. Search and its optimization. American Mathematical Monthly, 86(7):527–540, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goodman:1979:MOS

[102] Nicolas D. Goodman. Mathematics as an objective science. American Mathematical Monthly, 86(7):540–551, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Snapper:1979:WM

[103] Ernst Snapper. What is mathematics. American Mathematical Monthly, 86(7): 551–557, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cooke:1979:CLT

[104] R. L. Cooke. The Cantor-Lebesgue theorem. American Mathematical Monthly, 86(7):558–565, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ruckle:1979:BCT

[105] William H. Ruckle. The bounded consistency theorem. American Mathematical Monthly, 86(7):566–571, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eisenberg:1979:PHB

[106] Murray Eisenberg and Robert Guy. A proof of the hairy ball theorem. American Mathematical Monthly, 86(7):572– 574, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ginsburg:1979:MIT

[107] John Ginsburg and Bill Sands. Minimal infinite topological spaces. American Mathematical Monthly, 86(7):574– 576, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Haruki:1979:PQP

[108] Shigeru Haruki. A property of quadratic polynomials. American Mathematical Monthly, 86(7):577–579, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wagstaff:1979:SQA

[109] Samuel S. Wagstaff, Jr. Some questions about arithmetic progressions. American Mathematical Monthly, 86(7):579– 582, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rosenholtz:1979:TND

[110] Ira Rosenholtz. There is no differentiable metric for R^{n}. American Mathematical Monthly, 86(7):585–586, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rosenfeld:1979:DT

[111] Azriel Rosenfeld. Digital topology. American Mathematical Monthly, 86(8): 621–630, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pratt:1979:FHM

[112] John W. Pratt. Finding how many roots a polynomial has in (0, 1) or (0, ∞). American Mathematical Monthly, 86(8): 630–637, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Johnsonbaugh:1979:SAS

[113] Richard Johnsonbaugh. Summing an alternating series. American Mathematical Monthly, 86(8):637–648, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grosswald:1979:RAS

[114] Emil Grosswald. Recent applications of some old work of Laguerre. American Mathematical Monthly, 86(8):648– 658, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hrbacek:1979:NST

[115] Karel Hrbacek. Nonstandard set theory. American Mathematical Monthly, 86(8):659–677, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nurcombe:1979:SCT

[116] J. R. Nurcombe. A sequence of convergence tests. American Mathematical Monthly, 86(8):679–681, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bagby:1979:TPD

[117] Richard J. Bagby. Taylor polynomials and difference quotients. American Mathematical Monthly, 86(8):681– 684, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Minda:1979:CAC

[118] C. David Minda. Conformal automorphisms of circular regions. American Mathematical Monthly, 86(8):684– 686, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jordan:1979:TIS

[119] J. H. Jordan, Ray Walch, and R. J. Wisner. Triangles with integer sides. American Mathematical Monthly, 86(8):686– 689, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leighton:1979:ULT

[120] Walter Leighton. A useful lemma in the theory of second-order linear differential equations. American Mathematical Monthly, 86(8):693–694, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Miel:1979:AC

[121] George Miel. An algorithm for the calculation of π. American Mathematical Monthly, 86(8):694–697, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1979:PES

[122] D. H. Lehmer and Emma Lehmer. Picturesque exponential sums. I. American Mathematical Monthly, 86(9):725– 733, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mills:1979:DTC

[123] Wendell Mills, Boris Weisfeiler, and Allan M. Krall. Discovering theorems with a computer: the case of y′ = sin(xy). American Mathematical Monthly, 86(9): 733–739, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dawson:1979:GIT

[124] John Dawson. The Gödel incompleteness theorem from a length-of-proof perspective. American Mathematical Monthly, 86(9):740–747, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Greenberg:1979:ENE

[125] Marvin Jay Greenberg. Euclidean and non-Euclidean geometries without continuity. American Mathematical Monthly, 86(9):757–764, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schoenberg:1979:PSL

[126] I. J. Schoenberg. On a problem of Steinhaus on lattice points. American Mathematical Monthly, 86(9):765–766, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brown:1979:BUT

[127] A. L. Brown. The Borsuk-Ulam theorem and orthogonality in normed linear spaces. American Mathematical Monthly, 86(9):766–767, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Heiberg:1979:EFS

[128] C. H. Heiberg. Extrema for functions of several variables. American Mathematical Monthly, 86(9):767–770, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cassidy:1979:PCA

[129] Phyllis Joan Cassidy. Products of commutators are not always commutators: an example. American Mathematical Monthly, 86(9):772, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fearnley-Sander:1979:HGC

[130] Desmond Fearnley-Sander. Hermann Grassmann and the creation of linear algebra. American Mathematical Monthly, 86(10):809–817, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gale:1979:GHB

[131] David Gale. The game of Hex and the Brouwer fixed-point theorem. American Mathematical Monthly, 86(10):818– 827, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Greenberg:1979:RPP

[132] Marvin Jay Greenberg. The “real” projective plane without continuity. American Mathematical Monthly, 86(10):827– 832, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nijenhuis:1979:MPA

[133] Albert Nijenhuis. A minimal-path algorithm for the “money changing problem”. American Mathematical Monthly, 86(10):832–835, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Haimo:1979:CC

[134] Deborah Tepper Haimo and Franklin Tepper Haimo. Comments and complements. American Mathematical Monthly, 86(10):836–841, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kammler:1979:LCP

[135] David W. Kammler and Robert J. McGlinn. Local conditioning of parametric forms used to approximate continuous functions. American Mathematical Monthly, 86(10):841–845, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nelson:1979:PQ

[136] R. D. Nelson. Perfect quadrilaterals. American Mathematical Monthly, 86 (10):845–847, ???? 1979. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Briggs:1979:MC

[137] James M. Briggs and Thomas Schaffter.

Measure and cardinality. American Mathematical Monthly, 86(10):852–855, ???? 1979. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ecker:1980:CNC

[138] A. Ecker. Comment on the note:

“The congruence a^{r+s} ≡ a^{r} (mod m)” [Amer. Math. Monthly 85 (1978), no. 2, 97–100, MR 57 #12355] by A. E. Livingston and M. L. Livingston. American Mathematical Monthly, 87(10):811– 814, ???? 1980. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [48].

Wilder:1980:CAE

[139] R. L. Wilder. Correction and addendum to: “Evolution of the topological concept of ‘connected’ ” [Amer. Math. Monthly 85 (1978), no. 9, 720–726, MR 80f:01019a]. American Mathematical Monthly, 87(1):31–32, ???? 1980. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [34].

Golomb:1981:CCP

[140] Solomon W. Golomb. Corrections to: “Cyclotomic polynomials and factorization theorems” [Amer. Math. Monthly 85 (1978), no. 9, 734–737, MR 80e:10011]. American Mathematical Monthly, 88(5):338–339, ???? 1981. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [36].

McLain:1981:NWG

[141] Katherine E. McLain and Hugh M.

Edgar. A note on S. W. Golomb’s: “Cyclotomic polynomials and factorization theorems” [Amer. Math. Monthly 85 (1978), no. 9, 734–737, MR 80e:10011]. American Mathematical Monthly, 88 (10):753, ???? 1981. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [36].